Grade 10Algebra

What is the identity for (a+b+c+d)^2? Can u also give the explanation of derivation of this identity??

vaishnavi
10 Years agoGrade 10

1 Answer

Vikas TU
10 Years ago
identity for (a+b+c+d)^2 = 
 
a^2 + b^2 + c^2 + d^2 + 2(a(b+c+d) +b(c+d) + cd)
More identitites apart from this --->
a^3 + b^3 + c^3 = 3abc if a+b+c=0

a^4 + a^2 + 1 = (a^2+1-a) (a^2 +1 +a)
 


 
 
Remember it!